OpenStudy (anonymous):

v = sin (2x + 3y) what is dv/dy and dv/dx . ty

OpenStudy (anonymous):

\[f(x,y)=\sin(2x+3y),\]\[f_x(x,y)=2\cos(2x+3y),\]\[f_y(x,y)=3\cos(2x+3y).\]

OpenStudy (anonymous):

thanks alot